# CATP: Contextually Adaptive Token Pruning for Efficient and Enhanced Multimodal In‑Context Learning

### v1

@@ Abstract
Modern large vision-language models (LVLMs) convert each input image into a large set of tokens, far outnumbering the text tokens. Although this improves visual perception, it introduces severe image token redundancy. Because image tokens carry sparse information, many add little to reasoning, yet greatly increase inference cost. The emerging image token pruning methods tackle this issue by identifying the most important tokens and discarding the rest. These methods can raise efficiency with only modest performance loss. However, most of them only consider single-image tasks and overlook multimodal in-context learning (ICL), where redundancy is greater and efficiency is more critical. Redundant tokens weaken the advantage of multimodal ICL for rapid domain adaptation and cause unstable performance. Applying existing pruning methods in this setting leads to large accuracy drops, exposing a clear gap and the need for new techniques. Thus, we propose Contextually Adaptive Token Pruning (CATP), a training-free pruning method targeted at multimodal ICL. CATP consists of two stages that perform progressive pruning to fully account for the complex cross-modal interactions in the input sequence. After removing 77.8% of the image tokens, CATP produces an average performance gain of 0.6% over the vanilla model on four LVLMs and eight benchmarks, exceeding all baselines remarkably. Meanwhile, it effectively improves efficiency by achieving an average reduction of 10.78% in inference latency. CATP enhances the practical value of multimodal ICL and lays the groundwork for future progress in interleaved image-text scenarios.
@@ Introduction
Figure 1: (a) Example of a 3-shot in-context sequence. (b–d) show the effects of three image token pruning methods on this sequence. Attention-based methods tend to keep tokens at the bottom of the image due to attention sinks in interleaved inputs. Diversity-based methods lack semantic guidance from the entire context. In contrast, our proposed method CATP accounts for the complex cross-modal interactions within the sequence, leading to superior results.
@@
Large‑vision language models (LVLMs) have recently demonstrated impressive capabilities across a broad spectrum of visual‑language tasks (Liu et al. 2023a; Chen et al. 2024d; Wang et al. 2024). To equip LVLMs with novel‑domain capabilities while cutting the high cost of multimodal data preparation and training, multimodal in‑context learning has become an important extension (Doveh et al. 2024; Chen et al. 2024c). It allows a model to adapt on‑the‑fly by adding a small set of exemplars, known as in-context demonstrations (ICDs), directly into the prompt. This paradigm offers an elegant path toward task generalization without any parameter updates, making it increasingly popular for real‑world deployment (Li et al. 2025a).
@@
Although multimodal ICL offers a user-friendly and rapid solution, its input characteristics often impede the desired effects. As shown in Figure 1(a), every ICD and the subsequent query sample include an image, so the image token redundancy that is already a bottleneck in single-image tasks becomes even more acute. For instance, LLaVA-Next (Liu et al. 2024) converts each image into 576 tokens. On VizWiz (Gurari et al. 2018), a 2-shot ICL run requires 3.2 times the computation of single-image inference and an impressive 14.3 times that of text-only inference. The resulting latency and memory footprint conflict with the efficiency users expect. Moreover, the interleaved format introduces additional issues. Multimodal ICL involves several kinds of cross-modal interactions, and many uninformative tokens can complicate this process (Baldassini et al. 2024; Li et al. 2025c). They prevent the model from focusing on the visual regions in each image that match the paired text. They also make it harder, based on the query sample, to locate the critical reference information provided by the user.
@@
To ease this serious redundancy, training‑free image token pruning has become a promising direction. Its core idea is to define an importance criterion that measures the contribution of each image token during inference (Kim et al. 2022; Wen et al. 2025a). By retaining only the most informative tokens and removing or merging the rest, latency is effectively reduced with minimal accuracy loss. Existing pruning strategies fall mainly into two categories. (1) Diversity‑based methods operate after the vision encoder and projector but before the decoder, when image tokens have been projected into the decoder’s representation space yet have not interacted with texts (Wen et al. 2025b; Jeddi et al. 2025). These methods measure feature similarity among image tokens to penalize redundancy and encourage diversity, as exemplified by DivPrune (Alvar et al. 2025). (2) Attention‑based methods act inside the LLM decoder, where visual and textual embeddings are processed together (Zhang et al. 2024a; Arif et al. 2025; Xu et al. 2025). They estimate the importance of each image token by the amount of attention it receives from other tokens, as exemplified by FastV (Chen et al. 2024b).
@@
Despite being able to reduce up to 90% of the image tokens with a <<5% drop in performance, the above two pruning methods are developed and evaluated only in single‑image tasks and thus ignore the much more demanding multi‑image scenarios, especially interleaved image–text ICL. When we adapt FastV and DivPrune to multimodal ICL, their shortcomings become clear, as shown in the visualization of Figure 1 and the results on LLaVA‑Next‑7B in Figure 2. Both methods incur greater performance degradation in 4‑shot ICL compared to the single-image setting, and in many cases even underperform random pruning. This highlights a non‑negligible gap between single‑image tasks and multimodal ICL. To investigate this gap, we conduct further ICL experiments with both pruning methods under varied setups and obtain two key findings. First, in multimodal ICL, diversity-based methods fail to perform fine-grained pruning as they lack information from other images and texts, and thus cannot capture diverse cross-modal interactions. Second, while layer choice matters for attention‑based pruning, attention shifts (Zhang et al. 2025a) in every image bias all its variants toward text‑adjacent tokens, and the bias compounds to reduce accuracy. To this end, we question that
@@
To approach this target, we propose Contextually Adaptive Token Pruning (CATP), a training-free image token pruning method designed for multimodal ICL. CATP removes a ratio $R$ of the total image tokens in an in-context sequence. Based on our analysis of existing methods, CATP is divided into two stages. Stage I is applied between the projector and the decoder, and is tailored to the characteristics of multimodal ICL. It evaluates the token importance in two dimensions. The first is semantic alignment with the paired text, which ensures that key visual regions are retained as a foundation for effective ICL. The second is feature diversity, which prevents the tokens that matter more to the entire sequence from being discarded. Stage 2 acts on the shallow decoder layers. It adopts a novel progressive adaptation strategy that first treats all ICD image tokens as an integral context. By combining semantic relevance with layer-wise attention differences, it precisely identifies the most important context tokens. In the following layer, it prunes the query image tokens based on their semantic relevance to this distilled context, ensuring effective ICL after pruning. Extensive experiments demonstrate the effectiveness and generalization of CATP in multimodal ICL. It can improve both efficiency and performance, showing unique advantages.
@@
Figure 3: Average relative performance of pruning methods in 4-shot ICL: (a) showing diversity‑based pruning at 50%–90% ratios; (b–c) showing three different attention‑based pruning methods applied to various layers at 80% and 90% ratios.
Large Vision-language Models (LVLMs). The rapid rise of large language models (LLMs) gives birth to LVLMs (Li et al. 2024b, a; Chen et al. 2024d; Wang et al. 2024). A typical LVLM comprises three key modules: a vision encoder, a projector, and an LLM decoder. Although this architecture is generally effective, it has been found to be inefficient (Liu et al. 2023b; Li et al. 2024c). Input images usually occupy a large number of tokens, and these tokens carry much sparser information than text tokens, creating severe image token redundancy (Hu et al. 2024; Xing et al. 2024). For example, LLaVA‑1.5 (Li et al. 2024b) converts each image into 576 tokens, while LLaVA‑Next (Liu et al. 2024) and Qwen2.5-VL (Bai et al. 2025) increase the count up to 2880 tokens per image to handle high‑resolution inputs.
In-context Learning (ICL). ICL refers to a model’s ability to tackle new tasks by conditioning on a sequence of in-context demonstrations (ICDs) without updating any parameter (Brown et al. 2020; Dong et al. 2024; Wei et al. 2023). Due to its convenience, ICL has been widely adopted in NLP and extended to the multimodal domain. Current‑generation LVLMs, including the InternVL and QwenVL series, have already treated ICL as an essential capability, underscoring its high practical value (Chen et al. 2024d; Wang et al. 2024). However, the ideal efficiency of ICL is often diminished by image token redundancy, and the complex interactions in multimodal ICL make this issue even harder to resolve.
Image Token Pruning. Researchers have begun to tackle token redundancy by image token pruning. FastV (Chen et al. 2024b) first proposes to prune image tokens based on the attention they obtain in an early decoder layer. Following FastV, more attention‑based methods have appeared (Zhang et al. 2024a; Arif et al. 2025). VTW (Lin et al. 2025) proposes a token withdrawal strategy, removing all image tokens after a chosen layer. FitPrune (Ye et al. 2025) further leverages both cross‑attention and self‑attention of image tokens in every layer to measure their importance. Meanwhile, diversity-based methods argue that semantic diversity among image tokens is a better importance criterion than attention scores (Wen et al. 2025b; Jeddi et al. 2025). DivPrune (Alvar et al. 2025) performs pruning before image tokens enter the decoder and frames it as a Max-Min Diversity Problem. Although all these methods are effective, they were designed for single-image tasks and overlook multimodal ICL. We observe that their known limitations (Wen et al. 2025a; Zhang et al. 2025a) are amplified in this setting.
@@
where $D$ denotes the dimension of the decoder $f_{\phi}^{\theta}$ and $S^{I}_{i}$ is the number of tokens produced from each image. This number varies across LVLMs and can usually be set by the user. Diversity-based pruning methods are typically applied at this stage to reduce $S^{I}_{i}$. Meanwhile, each $i$-th text segment is converted into a set of text token embeddings $X^{T}_{i}\in\mathbb{R}^{S^{T}_{i}\times D}$ by $f_{\phi}^{\theta}$. After inserting the image tokens at their corresponding positions, we obtain the input token sequence:
@@
where $X_{sys}$ denotes the system prompt and the query sample is indexed as nn+11 for unified notation. The sequence $X$ is then fed into $f_{\phi}^{\theta}$, which performs an $N$-layer decoder forward pass to produce the final answer. Each layer contains a multihead attention module (MHA) and a feedforward MLP. The $l$-th layer maps the hidden states of $X_{ICS}$ to queries $Q_{l}$, keys $K_{l}$, and values $V_{l}$ by linear transformations, and the attention mechanism can be represented as:
$\mathbf{A}_{l}={softmax}(\frac{Q_{l}\,(K_{l})^{\top}+M}{\sqrt{D}})V_{l}\in\mathbb{R}^{S\times S}$ (4)
@@
where $S$ is the length of $X$ and $M$ is the causal mask. $K_{l}$ and $V_{l}$ are stored in the KV-cache for the decoding stage. Attention-based pruning methods are applied here to leverage $\mathbf{A}_{l}$ to discard less informative image tokens, preventing them from entering later layers. Both methods ultimately aim to identify the image tokens whose removal minimizes adverse effects on $\theta$’s decoding state given a pruning ratio.
@@
To further examine the gap revealed in Figure 2, we conduct additional 4-shot ICL evaluations on the six benchmarks using LLaVA-NeXT-7B and report the average relative performance (Avg. Rel) compared with the vanilla model. We first adapt DivPrune to ICL with pruning ratios ranging from 50% to 90%. As Figure 3(a) shows, it performs well at relatively low ratios and at 50% it even outperforms the vanilla model. When the ratio increases from 60% to 70%, its performance drops sharply and continues to decline. These results indicate that diversity-based methods fail to deliver fine-grained pruning in multimodal ICL. In these methods, each image token interacts only with tokens from the same image, so no information from other images or texts is considered. Meanwhile, multimodal ICL involves much richer cross‑modal interactions than single-image tasks, making this limitation more pronounced. Nonetheless, they remain useful as an initial filter before the decoder and facilitate further pruning.
@@
Next, we evaluate FastV by pruning 80% and 90% of image tokens at different decoder layers ranging from 2 to 16. FastV ranks tokens by the total attention they receive from all other tokens in a chosen layer. Considering the property of multimodal ICL, we introduce two alternatives: intra‑cross measures the attention that each image token receives from all text tokens within its own image-text pair, and query‑cross measures the attention that each image token receives from all tokens in the query sample. Intra‑cross performs best in very shallow layers, such as layer 2, where the model is focused on feature perception and alignment. As the layer index increases, the effectiveness of intra‑cross declines while FastV becomes superior. Query‑cross shows a sharp rise in the shallow layers, roughly layers 7 to 10, indicating that after feature perception, the LVLM shifts to query sample‑guided reasoning.
@@
However, none of the three methods reaches the desired performance in ICL. We attribute this to attention shifts that cause the model to focus excessively on image tokens positioned closer to the text segments (Zhang et al. 2025a). In interleaved settings, these shifts accumulate and make the attention signal even less indicative of true importance, as illustrated in Figure 1(b) and Appendix 2. Hence, statically relying on the attention assigned to image tokens is suboptimal for pruning in multimodal ICL.
@@
In multimodal ICL, each image is paired with text, so semantically aligned tokens are crucial for reasoning. However, the whole ICL process can depend more on features that interact with other ICDs or the query sample, such as complementary cross-sample cues. Thus, we measure the importance of this stage in two dimensions. The first dimension assesses the alignment between each image token and its paired text. The advantage of considering alignment at this stage is two-fold. After reducing the number of tokens, it promotes correct attention in the decoder by easing attention shifts (Li et al. 2025b). It also captures a key interaction in multimodal ICL that occurs within each image-text pair. This interaction is hard to isolate in the decoder as the forward pass merges semantics. The second assesses the diversity of the information carried by the image tokens. This keeps the retained tokens close to the feature distribution of the complete set and prevents tokens that contribute to the entire sequence from being removed. This matches the diversity-based methods, which are found to be suitable for initial filtering. To unify the two dimensions, we frame the task as a weighted submodular maximization problem.
@@
$Y_{i}^{*}=\arg\max_{\begin{subarray}{c}Y_{i}\subseteq X_{i}^{I},\\ |Y_{i}|=S_{i}^{I}\cdot(1-(R/2))\end{subarray}}\big{(}\mathcal{F}_{\text{div}}(Y_{i})+\lambda_{1}\cdot\mathcal{F}_{\text{align}}(Y_{i})\big{)},$ (5)
where $\lambda_{1}$ is a hyperparameter that balances the two objectives. The alignment function $\mathcal{F}_{\text{align}}(Y_{i})$ quantifies how well the selected image tokens align with the semantics of their paired text segment $X_{i}^{T}$. We obtain a single vector representing this semantics, $\bar{\mathbf{v}}_{i}$, by averaging all embeddings in $X_{i}^{T}$. Then, we compute the cosine similarity between each image token $x_{i,j}\in X^{I}_{i}$ and $\bar{\mathbf{v}}_{i}$ as $\mathcal{F}_{\text{align}}(Y_{i})$:
$\mathcal{F}_{\text{align}}(Y_{i})=\sum_{x_{i,j}\in Y_{i}}\text{sim}(x_{i,j},\bar{\mathbf{v}}_{i}).$ (6)
The diversity function $\mathcal{F}_{\text{div}}(Y_{i})$ is modeled with a facility location objective that drives the tokens in $Y_{i}$ to represent as diverse a set of features as possible. This function class is submodular and monotone, measuring the total similarity between all tokens in the ground set $X^{I}_{i}$ and their closest token in the chosen subset $Y_{i}$:
$\mathcal{F}_{\text{div}}(Y_{i})=\sum_{x\in X^{I}_{i}}\max_{y\in Y_{i}}\text{sim}(y,x).$ (7)
Since $\mathcal{F}_{\text{div}}(Y_{i})$ is submodular and $\mathcal{F}_{\text{align}}(Y_{i})$ is a modular function, their linear combination in Eq. 7 remains submodular. This property allows us to efficiently find a near-optimal solution using a standard greedy algorithm, making this stage a lightweight and powerful filter.
@@
We set the starting layer index of progressive adaptation to $K$. At this layer, we treat all ICDs as an integral context and prune their image tokens altogether until the overall ratio $R$ is reached. Thus, unlike existing methods, the pruning ratio inside each ICD is not fixed. It adapts to the importance of each ICD to the context as a whole. Because not every ICD contributes equally to ICL and some are even detrimental, this strategy preserves post-pruning performance. We use the query sample as a lens to prune the context. First, we compute a focused query vector $\bar{\mathbf{v}}_{\text{q}}^{K}$ by applying average pooling to the hidden states of all tokens in the query sample at layer $K$. We then score each context image token $c$ by comprehensively considering its relevance to the query and its dynamic importance growth. This growth is measured by the change in attention it receives from the query sample’s last token $t_{ql}$ between layers:
@@
where $\mathbf{A}$ is the attention weight matrix defined in Eq. 4. $idx_{ql}$ and $idx_{c}$ denote the token index of $t_{ql}$ and $c$, respectively. The full importance score is defined as:
$\mathcal{S}_{\text{context}}(c)=\Delta\mathcal{A}(c)+\lambda_{2}\cdot\mathrm{sim}(\mathbf{h}_{c}^{K},\bar{\mathbf{v}}_{\text{q}}^{K})$ (9)
where $\mathbf{h}_{c}^{K}$ denotes the hidden states of $c$ at layer $K$ and $\lambda_{2}$ is a hyperparameter. We globally rank all context image tokens by this score and prune them in ascending order. This step simultaneously distills the large context, allowing the LVLM to conduct more targeted and in-depth reasoning. Next, at layer $K+1$ we perform context-guided pruning on the query sample’s image tokens. We apply average pooling to the hidden states of the pruned context at this layer to obtain a distilled context vector $\bar{\mathbf{v}}_{c}^{K+1}$. To maintain a direct and focused mechanism, the importance of each query image token $q$ is only determined by its relevance to the distilled context:
@@
where $\mathbf{h}_{q}^{K+1}$ is the hidden states of $q$ at layer $K+1$. After pruning the query sample’s image tokens with this score, CATP is complete. Only $R$ of the total image tokens in $X$ proceeds to layer $K+2$ and all subsequent inference.

### v2

@@ Abstract
Modern large vision-language models (LVLMs) convert each input image into a large set of tokens that far outnumber the text tokens. Although this improves visual perception, it also introduces severe image token redundancy. Because image tokens contain sparse information, many contribute little to reasoning but greatly increase inference cost. Recent image token pruning methods address this issue by identifying important tokens and removing the rest. These methods improve efficiency with only small performance drops. However, most of them focus on single-image tasks and overlook multimodal in-context learning (ICL), where redundancy is higher and efficiency is more important. Redundant tokens weaken the advantage of multimodal ICL for rapid domain adaptation and lead to unstable performance. When existing pruning methods are applied in this setting, they cause large accuracy drops, which exposes a clear gap and the need for new approaches. To address this, we propose Contextually Adaptive Token Pruning (CATP), a training-free pruning method designed for multimodal ICL. CATP uses two stages of progressive pruning that fully reflect the complex cross-modal interactions in the input sequence. After removing 77.8% of the image tokens, CATP achieves an average performance gain of 0.6% over the vanilla model on four LVLMs and eight benchmarks, clearly outperforming all baselines. At the same time, it improves efficiency by reducing inference latency by an average of 10.78%. CATP strengthens the practical value of multimodal ICL and lays the foundation for future progress in interleaved image-text settings.
@@ Introduction
Figure 1: (a) Example of a 3-shot in-context sequence. (b-d) show the effects of three image token pruning methods on this sequence. Attention-based methods tend to keep tokens at the bottom of the image due to attention sinks in interleaved inputs. Diversity-based methods lack semantic guidance from the entire context. In contrast, our proposed method CATP accounts for the complex cross-modal interactions within the sequence, leading to superior results.
@@
Large vision-language models (LVLMs) have recently shown strong performance across a wide range of visual-language tasks (vituning; intern1; qwen1). To equip LVLMs with new domain capabilities while reducing the high cost of multimodal data preparation and training, multimodal in-context learning has become an important extension (icl1; icl2). It allows a model to adapt on the fly by adding a small set of exemplars, known as in-context demonstrations (ICDs), directly into the prompt. This paradigm provides a simple path toward task generalization without any parameter updates, making it increasingly attractive for real-world use (icv2).
@@
Although multimodal ICL offers a user-friendly and rapid solution, its input characteristics often impede the desired effects. As shown in Figure 1(a), every ICD and the subsequent query sample include an image, so the image token redundancy that is already a bottleneck in single-image tasks becomes even more acute. For instance, LLaVA-Next (LLaVAnext) converts each image into 576 tokens. On VizWiz (VizWiz), a 2-shot ICL run requires 3.2 times the computation of single-image inference and an impressive 14.3 times that of text-only inference. The resulting latency and memory footprint conflict with the efficiency users expect. Moreover, the interleaved format introduces additional issues. Multimodal ICL involves several kinds of cross-modal interactions, and many uninformative tokens can complicate this process (pos; taco). They prevent the model from focusing on the visual regions in each image that match the paired text. They also make it harder to locate the key information provided by the user based on the query sample (vega; comm).
@@
To ease this serious redundancy, training‑free image token pruning has become a promising direction. Its core idea is to define an importance criterion that measures the contribution of each image token during inference (pruning1; pruning2). By retaining only the most informative tokens and removing or merging the rest, latency is effectively reduced with minimal accuracy loss. Existing pruning strategies fall mainly into two categories. (1) Diversity‑based methods operate after the vision encoder and projector but before the decoder, when image tokens have been projected into the decoder’s representation space yet have not interacted with texts (div1; div2). These methods measure feature similarity among image tokens to penalize redundancy and encourage diversity, as exemplified by DivPrune (div). (2) Attention‑based methods act inside the LLM decoder, where visual and textual embeddings are processed together (att1; att2; att3). They estimate the importance of each image token by the amount of attention it receives from other tokens, as exemplified by FastV (fastv).
@@
Despite being able to reduce up to 90% of the image tokens with less than a 5% drop in performance, the above two pruning methods are developed and evaluated only in single‑image tasks and thus ignore the much more demanding multi‑image scenarios, especially interleaved image-text ICL. When we adapt FastV and DivPrune to multimodal ICL, their shortcomings become clear, as shown in the visualization of Figure 1 and the results on LLaVA‑Next‑7B in Figure 2. Both methods incur greater performance degradation in 4‑shot ICL compared to the single-image setting, and in many cases even underperform random pruning. This highlights a non‑negligible gap between single‑image tasks and multimodal ICL. To investigate this gap, we conduct further ICL experiments with both pruning methods under varied setups and obtain two key findings. First, in multimodal ICL, diversity-based methods fail to perform fine-grained pruning as they lack information from other images and texts, and thus cannot capture diverse cross-modal interactions. Second, while layer choice matters for attention‑based pruning, attention shifts (limi1) in every image bias all its variants toward text‑adjacent tokens, and the bias compounds to reduce accuracy. To this end, we question that
@@
To approach this target, we propose Contextually Adaptive Token Pruning (CATP), a training-free image token pruning method designed for multimodal ICL. CATP removes a ratio $R$ of the total image tokens in an in-context sequence. Based on our analysis of existing methods, CATP is divided into two stages. Stage 1 is applied between the projector and the decoder, and is tailored to the characteristics of multimodal ICL. It evaluates the token importance in two dimensions. The first is semantic alignment with the paired text, which ensures that key visual regions are retained as a foundation for effective ICL. The second is feature diversity, which prevents the tokens that matter more to the entire sequence from being discarded. Stage 2 acts on the shallow decoder layers. It adopts a novel progressive adaptation strategy that first treats all ICD image tokens as an integral context. By combining semantic relevance with layer-wise attention differences, it precisely identifies the most important context tokens. In the following layer, it prunes the query image tokens based on their semantic relevance to this distilled context, ensuring effective ICL after pruning. Extensive experiments show that CATP is both effective and generalizable in multimodal ICL. It improves efficiency and performance at the same time and provides clear advantages over all existing baselines.
@@
Figure 3: Average relative performance of pruning methods in 4-shot ICL: (a) showing diversity‑based pruning at 50%-90% ratios; (b-c) showing three different attention‑based pruning methods applied to various layers at 80% and 90% ratios.
Large Vision-language Models (LVLMs). The rapid rise of large language models (LLMs) gives birth to LVLMs (llava; onevision; intern1; qwen1). A typical LVLM comprises three key modules: a vision encoder, a projector, and an LLM decoder. Although this architecture is generally effective, it has been found to be inefficient (token1; token2). Input images usually occupy a large number of tokens, and these tokens carry much sparser information than text tokens, creating severe image token redundancy (redun1; redun2). For example, LLaVA‑1.5 (llava) converts each image into 576 tokens, while LLaVA‑Next (LLaVAnext) and Qwen2.5-VL (qwen) increase the count up to several thousand tokens per image to handle high‑resolution inputs.
In-context Learning (ICL). ICL refers to a model’s ability to solve new tasks by conditioning on a sequence of in-context demonstrations (ICDs) without updating any parameters (iclr1; iclr2; iclr3). Because of its convenience, ICL has been widely used in NLP and extended to the multimodal setting. Current-generation LVLMs, including the InternVL and QwenVL series, already treat ICL as a core capability, which highlights its strong practical value (intern1; qwen1). However, the ideal efficiency of ICL is often reduced by image token redundancy, and the complex interactions in multimodal ICL make this problem even harder to address.
Image Token Pruning. Researchers have begun to tackle token redundancy by image token pruning. FastV (fastv) first proposes to prune image tokens based on the attention they obtain in an early decoder layer. Following FastV, more attention‑based methods have appeared (att1; att2). VTW (vtw) proposes a token withdrawal strategy, removing all image tokens after a chosen layer. FitPrune (fit) further leverages both cross‑attention and self‑attention of image tokens in every layer to measure their importance. Meanwhile, diversity-based methods argue that semantic diversity among image tokens is a better importance criterion than attention scores (div1; div2). DivPrune (div) performs pruning before image tokens enter the decoder and frames it as a Max-Min Diversity Problem. Although all these methods are effective, they were designed for single-image tasks and overlook multimodal ICL. We find that their known limitations (pruning2; limi1) become even more pronounced in this setting.
@@
where $D$ denotes the dimension of the LLM decoder $\phi^{\theta}$ and $S^{I}_{i}$ is the number of tokens produced from each image. This number varies across LVLMs and is usually user-configurable. Diversity-based pruning methods are typically applied at this stage to reduce $S^{I}_{i}$. Meanwhile, each $i$-th text segment is converted into a set of text token embeddings $X^{T}_{i}\in\mathbb{R}^{S^{T}_{i}\times D}$ by $\phi^{\theta}$. After inserting the image tokens at their corresponding positions, we obtain the input token sequence:
@@
where $X_{sys}$ denotes the system prompt and the query sample is indexed as nn+11 for unified notation. The sequence $X$ is then fed into $\phi^{\theta}$ and processed by an $N$-layer decoder to produce the final answer. Each layer contains a multi-head attention module (MHA) and a feedforward MLP. The $h$-th head in the $l$-th layer maps the hidden states of $X_{ICS}$ to queries $Q_{l,h}$, keys $K_{l,h}$, and values $V_{l,h}$ by linear transformations, and the head’s attention weight matrix can be represented as:
$\mathbf{A}_{l,h}={softmax}(\frac{Q_{l,h}\,K_{l,h}^{\top}+\mathbf{M}}{\sqrt{D_{k}}})\in\mathbb{R}^{S\times S}$ (4)
@@
where $S$ is the length of $X$ and $\mathbf{M}$ is the causal mask. $D_{k}$ is the head dimension of $\phi^{\theta}$. $K_{l,h}$ and $V_{l,h}$ are stored in the KV cache during the prefilling stage to prepare for subsequent decoding. Attention-based pruning methods are applied here to leverage $\mathbf{A}_{l,h}$ to discard less informative image tokens, preventing them from entering later layers. Both methods aim to identify the tokens whose removal minimizes the adverse effects on $\theta$’s decoding, given a specific pruning ratio.
@@
To further examine the gap revealed in Figure 2, we conduct additional 4-shot ICL evaluations on the six benchmarks using LLaVA-Next-7B and report the average relative performance (Avg. Rel) compared with the vanilla model. We first adapt DivPrune to ICL with pruning ratios ranging from 50% to 90%. As Figure 3(a) shows, it performs well at relatively low ratios and at 50% it even outperforms the vanilla model. When the ratio increases from 60% to 70%, its performance decreases sharply and continues to decline. These results indicate that diversity-based methods fail to deliver fine-grained pruning in multimodal ICL. In these methods, each image token interacts only with tokens from the same image, so no information from other images or texts is considered. Meanwhile, multimodal ICL involves much richer cross‑modal interactions than single-image tasks, making this limitation more pronounced. Nonetheless, they remain useful as an initial filter before the decoder.
@@
Next, we evaluate FastV by pruning 80% and 90% of image tokens at different decoder layers ranging from 2 to 16. FastV ranks tokens by the total attention they receive from all other tokens in a chosen layer. Considering the property of multimodal ICL, we introduce two alternatives: intra‑cross measures the attention that each image token receives from all text tokens within its own image-text pair, and query‑cross measures the attention that each image token receives from all tokens in the query sample. Intra‑cross performs best in very shallow layers, such as layer 2, where the model is focused on feature perception and alignment. As the layer index increases, the effectiveness of intra‑cross declines while FastV becomes superior. Query‑cross shows a sharp rise in the shallow layers, roughly layers 7 to 10, indicating that after perception, the LVLM shifts to query‑guided reasoning.
@@
However, none of the three methods reaches the desired performance in ICL. We attribute this to attention shifts that cause the model to focus excessively on image tokens positioned closer to the text segments (limi1). In interleaved settings, these shifts accumulate and make the attention signal even less indicative of true importance, as illustrated in Figure 1(b) and Appendix 2. Hence, statically relying on the attention assigned to image tokens is suboptimal for pruning in multimodal ICL.
@@
In multimodal ICL, each image is paired with text, so semantically aligned tokens are crucial for reasoning. However, the whole ICL process can depend more on features that interact with other ICDs or the query sample, such as complementary cross-sample cues. Thus, we measure the importance of this stage in two dimensions. The first dimension assesses the alignment between each image token and its paired text. The advantage of considering alignment at this stage is two-fold. After reducing the number of tokens, it promotes correct attention in the decoder by easing attention shifts (cama). It also captures a key interaction in multimodal ICL that occurs within each image-text pair. This interaction is hard to isolate in the decoder as the forward pass merges semantics. The second assesses the diversity of the information carried by the image tokens. This keeps the retained tokens close to the feature distribution of the complete set and prevents tokens that contribute to the entire sequence from being removed. This matches the diversity-based methods, which are found to be suitable for initial filtering. To unify the two dimensions, we frame the task as a weighted submodular maximization problem.
@@
$Y_{i}^{*}=\arg\max_{\begin{subarray}{c}Y_{i}\subseteq X_{i}^{I},\\ |Y_{i}|=\left\lfloor S_{i}^{I}\cdot(1-(R/2))\right\rfloor\end{subarray}}\big(\mathcal{F}_{\text{div}}(Y_{i})+\lambda_{1}\cdot\mathcal{F}_{\text{align}}(Y_{i})\big),$ (5)
where $\lambda_{1}$ is a hyperparameter that balances the two objectives. The alignment function $\mathcal{F}_{\text{align}}(Y_{i})$ quantifies how well the selected image tokens align with the semantics of their paired text segment $X_{i}^{T}$. We first compute a single vector representing the $i$-th text segment’s semantics, $\bar{\mathbf{v}}_{i}$, by averaging all embeddings in $X_{i}^{T}$. Then, we obtain $\mathcal{F}_{\text{align}}(Y_{i})$ using the similarity score between each image token $x_{i,j}\in X^{I}_{i}$ and $\bar{\mathbf{v}}_{i}$:
$\mathcal{F}_{\text{align}}(Y_{i})=\sum_{x_{i,j}\in Y_{i}}\hat{\text{sim}}(x_{i,j},\bar{\mathbf{v}}_{i}),$ (6)
[…]
$\mathcal{F}_{\text{div}}(Y_{i})=\sum_{x\in X^{I}_{i}}\max_{y\in Y_{i}}\hat{\text{sim}}(y,x).$ (7)
Since $\mathcal{F}_{\text{div}}(Y_{i})$ is submodular and $\mathcal{F}_{\text{align}}(Y_{i})$ is a modular function, their linear combination in Eq. 5 remains submodular. This property allows us to efficiently find a near-optimal pruning strategy using a standard greedy algorithm, making stage 1 a lightweight and effective pre-decoder filter.
@@
We set the starting layer index of progressive adaptation to $K$. At this layer, we treat all ICDs as an integral context and prune their image tokens altogether until the overall ratio $R$ is reached. Thus, unlike existing methods, the pruning ratio inside each ICD is not fixed. It adapts to the importance of each ICD to the context as a whole. Because not every ICD contributes equally to ICL and some are even detrimental, this strategy preserves post-pruning performance. We use the query sample as a lens to prune the context. First, we compute a focused query vector $\bar{\mathbf{v}}_{\text{q}}^{K}$ by applying average pooling to the hidden states of all tokens in the query sample at layer $K$. We then score each context image token $c$ by comprehensively considering its relevance to the query and its dynamic growth in importance This growth is measured by the change in attention it receives from the query sample’s last token $t_{ql}$ between layers:
@@
where $\mathbf{A}_{l}$ is the $l$-th layer’s attention weight matrix averaged over all heads. $idx_{ql}$ and $idx_{c}$ denote the token index of $t_{ql}$ and $c$, respectively. The full importance score is defined as:
$\mathcal{S}_{\text{context}}(c)=\Delta\mathcal{A}(c)+\lambda_{2}\cdot\mathrm{sim}(\mathbf{h}_{c}^{K},\bar{\mathbf{v}}_{\text{q}}^{K}),$ (9)
where $\mathbf{h}_{c}^{K}$ denotes the hidden states of $c$ at layer $K$ and $\lambda_{2}$ is a hyperparameter. We globally rank all context image tokens by this score and prune them in ascending order. This step simultaneously distills the large context, allowing the LVLM to conduct more targeted and in-depth reasoning. Next, at layer $K+1$ we perform context-guided pruning on the query sample’s image tokens. We apply average pooling to the hidden states of the pruned context at this layer to obtain a distilled context vector $\bar{\mathbf{v}}_{c}^{K+1}$. To maintain a focused mechanism, the importance of each query image token $q$ is only determined by its relevance to the distilled context:
@@
where $\mathbf{h}_{q}^{K+1}$ is the hidden states of $q$ at layer $K+1$. After pruning the query sample’s image tokens with this score, CATP is complete. Only 1R1-R of the total image tokens in $X$ proceeds to layer $K+2$ and all subsequent inference.